OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x

OpenStudy (anonymous):

Sum and Difference Formulas sin(A+B)=sin A cos B + cos A sin B sin(A-B)=sin A cos B - cos A sin B cos(A+B)=cos A cos B - sin A sin B cos(A-B)=cos A cos B + sin A sin B http://www.ies.co.jp/math/java/trig/kahote/kahote.html

OpenStudy (anonymous):

@Schrodinger spent alot of time covering this material with you yesterday. review that material.